# Worse than Random? An Embarrassingly Simple Probing Evaluation of Large Multimodal Models in Medical VQA

### v1

@@ A.3 Spine CT Scan
Random Choice acc 25 25 50 30.85 31.06
acc w.o. adv. pair 50 50 / 36.42 36.62
LLaVA-v1-7B acc 22.98 44.82 50 0 0
acc w.o. adv. pair 98.85 91.95 / 97.7 100
LLaVA-v1.6-7B acc 4.59 72.41 0 0 1.28
acc w.o. adv. pair 98.85 85.05 / 65.51 70.51
LLaVA-Med acc 2.29 11.49 50 11.49 6.41
acc w.o. adv. pair 94.25 28.73 / 26.43 79.48
MiniGPT-v2 acc 1.14 41.37 0 12.64 5.12
acc w.o. adv. pair 100 62.06 / 68.96 78.2
CheXagent acc 4.59 27.58 50 10.34 2.56
acc w.o. adv. pair 16.09 55.17 / 27.58 33.33
GPT-4V acc 81.39 69.76 0 33.73 25.97
acc w.o. adv. pair 81.39 81.39 / 39.53 42.85
Gemini Pro acc 87.2 77.9 50 22.09 25.97
acc w.o. adv. pair 94.18 93.02 / 65.11 50.64
@@ A.4 Abdominal CT Scan
Random Choice acc 25 25 50 35.53 37.03
acc w.o. adv. pair 50 50 / 40.86 41.88
LLaVA-v1-7B acc 26.49 54.19 50 0.53 0
acc w.o. adv. pair 100 100 / 94.54 99.83
LLaVA-v1.6-7B acc 1.86 82.82 41.42 1.06 0.66
acc w.o. adv. pair 99.86 96.4 / 70.17 87.76
LLaVA-Med acc 5.05 45 30 15.44 5.28
acc w.o. adv. pair 98 93.6 / 33.28 86.11
MiniGPT-v2 acc 0 37.15 48.57 6.12 2.14
acc w.o. adv. pair 99.2 91.34 / 83.62 90.74
CheXagent acc 25.03 38.21 52.85 15.57 6.61
acc w.o. adv. pair 56.99 80.29 / 51.26 55.53
GPT-4V acc 95.72 72.72 45.71 27 23.25
acc w.o. adv. pair 96.39 92.11 / 40.9 34.71
Gemini Pro acc 98.31 69.19 65.71 28.79 20.39
acc w.o. adv. pair 99.2 96.8 / 61.86 58.2
@@ A.5 Brain MRI
Random Choice acc 25 25 50 36.7 36.64
acc w.o. adv. pair 50 50 / 41.47 41.49
LLaVA-v1-7B acc 1.23 32.86 50 0.53 0
acc w.o. adv. pair 97.17 97.34 / 94.52 99.75
LLaVA-v1.6-7B acc 17.49 88.51 28.57 0.53 0.48
acc w.o. adv. pair 99.49 89.57 / 54.77 67.63
LLaVA-Med acc 3 8.12 23.21 14.66 2.91
acc w.o. adv. pair 90.45 44.34 / 22.08 89.29
MiniGPT-v2 acc 1.94 96.64 50 15.72 4.37
acc w.o. adv. pair 99.82 100 / 66.78 88.8
CheXagent acc 0 0 50 10.77 6.81
acc w.o. adv. pair 1.41 0 / 16.07 23.66
GPT-4V acc 96.99 94.33 58.92 36.1 27.8
acc w.o. adv. pair 97.52 95.39 / 54.69 43.17
Gemini Pro acc 95.22 94.87 78.57 35.51 19.7
acc w.o. adv. pair 96.28 99.29 / 56.89 37.71
@@ A.6 Chest MRI
Random Choice acc 25 25 50 34.18 34.11
acc w.o. adv. pair 50 50 / 39.84 39.88
LLaVA-v1-7B acc 0 35 50 0 0
acc w.o. adv. pair 80 92.5 / 90 100
LLaVA-v1.6-7B acc 5 32.5 37.5 0 0
acc w.o. adv. pair 90 50 / 50 76.47
LLaVA-Med acc 5 45 12.5 12.5 5.88
acc w.o. adv. pair 82.5 82.5 / 17.5 82.35
MiniGPT-v2 acc 0 35 50 10 8.82
acc w.o. adv. pair 92.5 62.5 / 70 79.41
CheXagent acc 0 75 50 15 0
acc w.o. adv. pair 5 100 / 27.5 23.52
GPT-4V acc 76.92 51.28 37.5 25.64 18.18
acc w.o. adv. pair 76.92 53.84 / 30.76 30.3
Gemini Pro acc 87.5 62.5 37.5 17.5 11.76
acc w.o. adv. pair 87.5 82.5 / 55 44.11
@@ A.7 Spine MRI
Random Choice acc 25 25 50 31.51 31.52
acc w.o. adv. pair 50 50 / 37.55 37.66
LLaVA-v1-7B acc 0 32.09 50 50 0.3
acc w.o. adv. pair 98.45 98.76 / 93.51 99.32
LLaVA-v1.6-7B acc 3.08 86.72 27.77 0.3 0
acc w.o. adv. pair 99.69 95.37 / 50.3 57.77
LLaVA-Med acc 1.54 5.24 36.11 12.96 5.4
acc w.o. adv. pair 88.27 23.45 / 17.96 77.7
MiniGPT-v2 acc 0.3 49.69 52.77 6.79 2.02
acc w.o. adv. pair 100 84.56 / 80.86 87.83
CheXagent acc 0 13.58 47.22 15.43 2.7
acc w.o. adv. pair 1.54 27.77 / 22.53 25.33
GPT-4V acc 96.28 90.71 55.55 22.6 15.59
acc w.o. adv. pair 96.28 95.66 / 32.19 26.44
Gemini Pro acc 98.13 88.81 57.14 24.53 14.91
acc w.o. adv. pair 98.13 97.82 / 50 39.32
@@ A.8 Abdominal MRI
Random Choice acc 25 25 50 37.13 38.26
acc w.o. adv. pair 50 50 / 41.81 42.23
LLaVA-v1-7B acc 0 39.28 50 2.38 0
acc w.o. adv. pair 96.42 100 / 90.47 100
LLaVA-v1.6-7B acc 2.38 73.8 35.71 1.19 0
acc w.o. adv. pair 98.8 92.85 / 63.09 80.3
LLaVA-Med acc 2.38 47.61 50 14.28 9.09
acc w.o. adv. pair 79.76 95.23 / 17.85 71.21
MiniGPT-v2 acc 0 36.9 50 8.33 4.54
acc w.o. adv. pair 98.8 91.66 / 82.14 93.93
CheXagent acc 0 26.19 50 11.9 10.6
acc w.o. adv. pair 0 59.52 / 30.95 30.3
GPT-4V acc 86.9 75 50 27.38 25.75
acc w.o. adv. pair 86.9 86.9 / 50 42.42
Gemini Pro acc 89.28 72.61 85.71 28.57 25.75
acc w.o. adv. pair 91.66 94.04 / 67.85 56.06
@@ A.9 Brain X-ray
Random Choice acc 25 25 50 44.77 47.08
acc w.o. adv. pair 50 50 / 46.51 48.33
LLaVA-v1-7B acc 45.56 26.58 50 0 0
acc w.o. adv. pair 100 72.15 / 96.2 100
LLaVA-v1.6-7B acc 11.39 13.92 16.66 8.86 4.44
acc w.o. adv. pair 94.93 20.25 / 72.15 91.11
LLaVA-Med acc 8.86 8.86 0 20.25 4.44
acc w.o. adv. pair 100 27.84 / 24.05 88.88
MiniGPT-v2 acc 18.98 83.54 50 18.98 17.77
acc w.o. adv. pair 98.73 88.6 / 63.29 73.33
CheXagent acc 84.81 0 50 12.65 8.88
acc w.o. adv. pair 88.6 1.26 / 21.51 22.22
GPT-4V acc 82.05 8.97 33.33 43.58 22.72
acc w.o. adv. pair 83.33 8.97 / 50 36.36
Gemini Pro acc 89.87 51.89 50 31.64 31.11
acc w.o. adv. pair 93.67 62.02 / 50.63 46.66
@@ A.10 Chest X-ray
Random Choice acc 25 25 50 37.59 37.08
acc w.o. adv. pair 50 50 / 42.15 41.87
LLaVA-v1-7B acc 28.75 36.57 50 0.12 0.11
acc w.o. adv. pair 99.9 99.83 / 68.52 99.96
LLaVA-v1.6-7B acc 7.11 83.97 47.94 5.89 1.52
acc w.o. adv. pair 99.51 97.53 / 53.94 94.02
LLaVA-Med acc 6.25 39.77 40.24 26.28 6.14
acc w.o. adv. pair 99 90.6 / 35.67 94.34
MiniGPT-v2 acc 4.93 94.07 50.05 18.78 11.94
acc w.o. adv. pair 99.96 99.39 / 55.19 67.49
CheXagent acc 53.68 39.64 76.59 42.75 8.9
acc w.o. adv. pair 62.59 47.53 / 58.1 36.93
GPT-4V acc 91.53 67.51 53.18 39.35 21.35
acc w.o. adv. pair 91.88 94.09 / 46.18 34.38
Gemini Pro acc 98.07 76.74 61.29 25.83 15.31
acc w.o. adv. pair 99.29 99.67 / 57.83 66.42
@@ A.11 Spine X-ray
Random Choice acc 25 25 50 30.95 30.99
acc w.o. adv. pair 50 50 / 37.54 37.51
LLaVA-v1-7B acc 44.55 45.04 50 0.49 0
acc w.o. adv. pair 98.51 98.51 / 93.56 98.85
LLaVA-v1.6-7B acc 4.45 82.67 33.33 1.48 0.57
acc w.o. adv. pair 99.5 95.54 / 61.38 77.58
LLaVA-Med acc 8.41 7.92 33.33 17.82 5.74
acc w.o. adv. pair 97.02 28.21 / 25.24 78.16
MiniGPT-v2 acc 2.97 52.47 58.33 16.33 4.02
acc w.o. adv. pair 99 75.24 / 61.38 69.54
CheXagent acc 82.17 20.29 62.5 16.83 0.57
acc w.o. adv. pair 90.09 40.59 / 30.69 31.6
GPT-4V acc 85.57 72.13 47.82 29.85 18.49
acc w.o. adv. pair 85.57 81.09 / 37.13 29.47
Gemini Pro acc 95.02 70.14 70.83 17.91 19.07
acc w.o. adv. pair 97.01 93.53 / 79.1 58.95

### v5

@@ A.3 Spine CT Scan
Random Choice Acc. with adv. pairs 25 25 50 30.85 31.06
LLaVA-v1 Acc. with adv. pairs 22.98 44.82 50 0 0
Avg. acc. 61.49 70.68 / 49.47 50.00
LLaVA-v1.6 Acc. with adv. pairs 4.59 72.41 0 0 1.28
Avg. acc. 52.29 83.90 / 37.66 41.07
MiniGPT-v2 Acc. with adv. pairs 1.14 41.37 0 12.64 5.12
Avg. acc. 50.57 58.62 / 54.41 51.21
LLaVA-Med-v1 Acc. with adv. pairs 2.29 11.49 50 11.49 6.41
Avg. acc. 48.27 30.45 / 46.37 48.77
LLaVA-Med-v1.5 Acc. with adv. pairs 32.18 67.81 50.0 9.19 14.10
Avg. acc. 65.51 83.33 / 55.23 51.27
BiomedGPT Acc. with adv. pairs 28.73 8.04 0 6.89 2.56
Avg. acc. 63.79 53.44 / 50.00 33.27
Med-Flamingo Acc. with adv. pairs 6.89 39.08 50 14.94 8.97
Avg. acc. 53.44 68.39 / 53.92 52.22
CheXagent Acc. with adv. pairs 4.59 27.58 50 10.34 2.56
Avg. acc. 34.48 58.04 / 49.45 50.20
GPT-4o Acc. with adv. pairs 87.35 76.74 0 30.23 20.77
Avg. acc. 93.10 88.37 / 66.01 60.08
GPT-4V Acc. with adv. pairs 81.39 69.76 0 33.73 25.97
Avg. acc. 89.53 84.30 / 65.77 63.13
Gemini Pro Acc. with adv. pairs 87.2 77.9 50 22.09 25.97
Avg. acc. 92.44 88.95 / 61.64 64.94
@@ A.4 Abdominal CT Scan
Random Choice Acc. with adv. pairs 25 25 50 35.53 37.03
LLaVA-v1 Acc. with adv. pairs 26.49 54.19 50 0.53 0
Avg. acc. 63.24 77.09 / 47.70 50.00
LLaVA-v1.6 Acc. with adv. pairs 1.86 82.82 41.42 1.06 0.66
Avg. acc. 50.93 91.07 / 38.36 45.82
MiniGPT-v2 Acc. with adv. pairs 0 37.15 48.57 6.12 2.14
Avg. acc. 50.00 66.97 / 48.49 50.22
LLaVA-Med-v1 Acc. with adv. pairs 5.05 45 30 15.44 5.28
Avg. acc. 51.53 70.90 / 45.13 49.24
LLaVA-Med-v1.5 Acc. with adv. pairs 51.93 67.64 48.57 11.31 16.03
Avg. acc. 75.96 83.42 / 52.86 65.61
BiomedGPT Acc. with adv. pairs 67.77 12.38 57.14 15.31 4.62
Avg. acc. 83.75 55.52 / 54.49 45.06
Med-Flamingo Acc. with adv. pairs 1.73 35.55 50 20.37 8.26
Avg. acc. 50.86 67.57 / 51.03 49.46
CheXagent Acc. with adv. pairs 25.03 38.21 52.85 15.57 6.61
Avg. acc. 51.46 65.71 / 51.08 50.19
GPT-4o Acc. with adv. pairs 97.99 65.28 51.42 23.12 28.23
Avg. acc. 98.93 81.50 / 58.24 64.59
GPT-4V Acc. with adv. pairs 95.72 72.72 45.71 27 23.25
Avg. acc. 97.72 85.56 / 58.92 60.02
Gemini Pro Acc. with adv. pairs 98.31 69.19 65.71 28.79 20.39
Avg. acc. 99.00 84.20 / 61.03 59.27
@@ A.5 Brain MRI
Random Choice Acc. with adv. pairs 25 25 50 36.7 36.64
LLaVA-v1 Acc. with adv. pairs 1.23 32.86 50 0.53 0
Avg. acc. 49.29 65.37 / 47.85 49.87
LLaVA-v1.6 Acc. with adv. pairs 17.49 88.51 28.57 0.53 0.48
Avg. acc. 58.74 93.10 / 31.73 37.46
MiniGPT-v2 Acc. with adv. pairs 1.94 96.64 50 15.72 4.37
Avg. acc. 50.88 98.32 / 52.16 50.51
LLaVA-Med-v1 Acc. with adv. pairs 3 8.12 23.21 14.66 2.91
Avg. acc. 47.08 32.50 / 47.72 48.35
LLaVA-Med-v1.5 Acc. with adv. pairs 75.61 84.98 42.85 13.78 13.62
Avg. acc. 87.80 92.40 / 53.37 53.52
BiomedGPT Acc. with adv. pairs 15.37 12.36 44.64 11.48 2.67
Avg. acc. 54.41 56.00 / 51.26 42.06
Med-Flamingo Acc. with adv. pairs 0.35 13.60 50 10.77 3.16
Avg. acc. 47.61 51.32 / 48.27 50.01
CheXagent Acc. with adv. pairs 0 0 50 10.77 6.81
Avg. acc. 20.40 21.99 / 50.37 51.87
GPT-4o Acc. with adv. pairs 97.69 97.34 66.07 25.84 30.24
Avg. acc. 98.58 98.67 / 61.05 66.13
GPT-4V Acc. with adv. pairs 96.99 94.33 58.92 36.1 27.8
Avg. acc. 98.40 97.07 / 65.89 62.38
Gemini Pro Acc. with adv. pairs 95.22 94.87 78.57 35.51 19.7
Avg. acc. 97.26 97.34 / 65.59 59.67
@@ A.6 Chest MRI
Random Choice Acc. with adv. pairs 25 25 50 34.18 34.11
LLaVA-v1 Acc. with adv. pairs 0 35 50 0 0
Avg. acc. 41.25 66.25 / 45.00 50.00
LLaVA-v1.6 Acc. with adv. pairs 5 32.5 37.5 0 0
Avg. acc. 51.24 56.25 / 31.35 43.01
MiniGPT-v2 Acc. with adv. pairs 0 35 50 10 8.82
Avg. acc. 47.50 62.50 / 47.91 49.50
LLaVA-Med-v1 Acc. with adv. pairs 5 45 12.5 12.5 5.88
Avg. acc. 43.75 68.75 / 49.06 46.32
LLaVA-Med-v1.5 Acc. with adv. pairs 50.00 35.00 50.00 12.5 11.76
Avg. acc. 72.5 62.5 / 53.75 53.92
BiomedGPT Acc. with adv. pairs 0.00 5.00 50.00 10.00 2.94
Avg. acc. 40.00 51.24 / 51.04 49.01
Med-Flamingo Acc. with adv. pairs 2.50 45.00 50 10.00 8.82
Avg. acc. 43.75 72.50 / 48.75 47.79
CheXagent Acc. with adv. pairs 0 75 50 15 0
Avg. acc. 17.50 87.50 / 44.58 47.05
GPT-4o Acc. with adv. pairs 90.00 35.89 62.50 17.94 24.24
Avg. acc. 93.75 65.38 / 54.80 61.36
GPT-4V Acc. with adv. pairs 76.92 51.28 37.5 25.64 18.18
Avg. acc. 86.25 71.79 / 58.11 61.74
Gemini Pro Acc. with adv. pairs 87.5 62.5 37.5 17.5 11.76
Avg. acc. 91.25 77.50 / 54.89 56.86
@@ A.7 Spine MRI
Random Choice Acc. with adv. pairs 25 25 50 31.51 31.52
LLaVA-v1 Acc. with adv. pairs 0 32.09 50 50 0.3
Avg. acc. 49.22 65.58 / 47.15 49.88
LLaVA-v1.6 Acc. with adv. pairs 3.08 86.72 27.77 0.30 0
Avg. acc. 51.54 92.74 / 30.59 34.37
MiniGPT-v2 Acc. with adv. pairs 0.3 49.69 52.77 6.79 2.02
Avg. acc. 50.15 70.52 / 48.97 50.01
LLaVA-Med-v1 Acc. with adv. pairs 1.54 5.24 36.11 12.96 5.4
Avg. acc. 45.06 24.07 / 48.52 48.04
LLaVA-Med-v1.5 Acc. with adv. pairs 70.67 84.56 50.00 11.11 11.14
Avg. acc. 84.72 91.97 / 52.40 51.89
BiomedGPT Acc. with adv. pairs 0.30 5.86 50.00 7.71 3.04
Avg. acc. 45.06 52.77 / 51.31 44.04
Med-Flamingo Acc. with adv. pairs 0.30 29.93 50 “ 17.90 5.40
Avg. acc. 50.00 64.50 / 50.54 50.14
CheXagent Acc. with adv. pairs 0 13.58 47.22 15.43 2.7
Avg. acc. 22.53 44.44 / 51.28 48.54
GPT-4o Acc. with adv. pairs 98.44 84.52 63.88 19.50 24.40
Avg. acc. 98.91 91.95 / 55.46 63.70
GPT-4V Acc. with adv. pairs 96.28 90.71 55.55 22.6 15.59
Avg. acc. 97.51 94.73 / 58.89 57.52
Gemini Pro Acc. with adv. pairs 98.13 88.81 57.14 24.53 14.91
Avg. acc. 98.75 94.09 / 59.19 58.20
@@ A.8 Abdominal MRI
Random Choice Acc. with adv. pairs 25 25 50 37.13 38.26
LLaVA-v1 Acc. with adv. pairs 0 39.28 50.00 2.38 0
Avg. acc. 48.22 69.64 / 46.42 50.00
LLaVA-v1.6 Acc. with adv. pairs 2.38 73.8 35.71 1.19 0
Avg. acc. 51.19 85.11 / 35.46 44.06
MiniGPT-v2 Acc. with adv. pairs 0 36.9 50 8.33 4.54
Avg. acc. 50.00 67.26 / 47.51 51.70
LLaVA-Med-v1 Acc. with adv. pairs 2.38 47.61 50.00 14.28 9.09
Avg. acc. 41.66 72.61 / 47.42 46.46
LLaVA-Med-v1.5 Acc. with adv. pairs 51.19 65.47 50.00 13.09 16.66
Avg. acc. 75.59 81.54 / 54.31 56.37
BiomedGPT Acc. with adv. pairs 1.19 3.57 50.00 14.28 1.51
Avg. acc. 38.69 50.00 / 51.33 46.46
Med-Flamingo Acc. with adv. pairs 2.38 27.38 50.00 20.23 3.03
Avg. acc. 50.59 62.50 / 49.55 50.50
CheXagent Acc. with adv. pairs 0 26.19 50.00 11.9 10.6
Avg. acc. 19.04 56.54 / 49.20 49.62
GPT-4o Acc. with adv. pairs 91.66 67.85 64.28 21.42 39.39
Avg. acc. 95.83 81.54 / 55.30 70.51
GPT-4V Acc. with adv. pairs 86.9 75 50 27.38 25.75
Avg. acc. 92.26 85.71 / 58.58 58.77
Gemini Pro Acc. with adv. pairs 89.28 72.61 85.71 28.57 25.75
Avg. acc. 94.04 86.30 / 63.39 60.98
@@ A.9 Brain X-ray
Random Choice Acc. with adv. pairs 25 25 50 44.77 47.08
LLaVA-v1 Acc. with adv. pairs 45.56 26.58 50 0 0
Avg. acc. 72.78 51.89 / 48.10 50.00
LLaVA-v1.6 Acc. with adv. pairs 11.39 13.92 16.66 8.86 4.44
Avg. acc. 55.06 48.10 / 45.04 48.88
MiniGPT-v2 Acc. with adv. pairs 18.98 83.54 50 18.98 17.77
Avg. acc. 59.49 89.87 / 51.37 52.22
LLaVA-Med-v1 Acc. with adv. pairs 8.86 8.86 0 20.25 4.44
Avg. acc. 54.43 31.01 / 51.16 48.33
LLaVA-Med-v1.5 Acc. with adv. pairs 49.36 31.64 50.00 8.86 13.33
Avg. acc. 73.41 56.96 / 53.16 55.55
BiomedGPT Acc. with adv. pairs 12.65 6.32 50.00 11.39 2.22
Avg. acc. 53.16 49.36 / 52.95 43.33
Med-Flamingo Acc. with adv. pairs 8.86 0 50 22.78 8.88
Avg. acc. 54.43 15.18 / 50.73 48.33
CheXagent Acc. with adv. pairs 84.81 0 50 12.65 8.88
Avg. acc. 92.40 29.74 / 51.16 55.00
GPT-4o Acc. with adv. pairs 94.93 52.56 66.66 37.17 40.90
Avg. acc. 96.20 73.71 / 62.07 69.31
GPT-4V Acc. with adv. pairs 82.05 8.97 33.33 43.58 22.72
Avg. acc. 90.38 47.43 / 68.48 59.09
Gemini Pro Acc. with adv. pairs 89.87 51.89 50 31.64 31.11
Avg. acc. 93.03 74.05 / 61.81 63.88
@@ A.10 Chest X-ray
Random Choice Acc. with adv. pairs 25 25 50 37.59 37.08
LLaVA-v1 Acc. with adv. pairs 28.75 36.57 50 0.12 0.11
Avg. acc. 64.37 68.25 / 34.41 50.05
LLaVA-v1.6 Acc. with adv. pairs 7.11 83.97 47.94 5.89 1.52
Avg. acc. 53.49 91.61 / 34.52 48.85
MiniGPT-v2 Acc. with adv. pairs 4.93 94.07 50.05 18.78 11.94
Avg. acc. 52.46 96.98 / 46.09 53.15
LLaVA-Med-v1 Acc. with adv. pairs 6.25 39.77 40.24 26.28 6.14
Avg. acc. 52.62 67.19 / 50.78 51.34
LLaVA-Med-v1.5 Acc. with adv. pairs 55.44 65.48 49.53 31.82 9.78
Avg. acc. 77.67 82.69 / 62.70 54.22
BiomedGPT Acc. with adv. pairs 91.34 86.05 50.00 16.92 9.08
Avg. acc. 95.46 92.93 / 43.00 41.46
Med-Flamingo Acc. with adv. pairs 80.92 90.00 50 35.83 5.24
Avg. acc. 90.46 95.00 / 63.47 48.00
CheXagent Acc. with adv. pairs 53.68 39.64 76.59 42.75 9.38
Avg. acc. 76.84 69.82 / 70.80 54.00
GPT-4o Acc. with adv. pairs 97.97 62.98 62.01 32.13 21.81
Avg. acc. 98.81 81.39 / 59.35 59.95
GPT-4V Acc. with adv. pairs 91.53 67.51 53.18 39.35 21.35
Avg. acc. 95.62 83.37 / 64.69 55.64
Gemini Pro Acc. with adv. pairs 98.07 76.74 61.29 25.83 15.31
Avg. acc. 98.94 88.32 / 52.22 54.97
@@ A.11 Spine X-ray
Random Choice Acc. with adv. pairs 25 25 50 30.95 30.99
LLaVA-v1 Acc. with adv. pairs 44.55 45.04 50 0.49 0
Avg. acc. 72.27 71.78 / 47.32 49.42
LLaVA-v1.6 Acc. with adv. pairs 4.45 82.67 33.33 1.48 0.57
Avg. acc. 52.22 90.84 / 35.87 42.02
MiniGPT-v2 Acc. with adv. pairs 2.97 52.47 58.33 16.33 4.02
Avg. acc. 51.48 71.78 / 53.84 51.07
LLaVA-Med-v1 Acc. with adv. pairs 8.41 7.92 33.33 17.82 5.74
Avg. acc. 52.72 28.96 / 52.82 47.58
LLaVA-Med-v1.5 Acc. with adv. pairs 46.53 71.78 50.00 14.85 13.32
Avg. acc. 73.01 85.89 / 55.78 54.79
BiomedGPT Acc. with adv. pairs 40.09 16.83 58.33 12.37 2.87
Avg. acc. 68.06 55.19 / 50.27 40.77
Med-Flamingo Acc. with adv. pairs 14.35 25.24 50 14.85 5.17
Avg. acc. 57.17 62.12 / 51.09 48.38
CheXagent Acc. with adv. pairs 82.17 20.29 62.5 16.83 0.57
Avg. acc. 91.08 50.74 / 52.70 48.70
GPT-4o Acc. with adv. pairs 95.54 79.70 47.82 34.15 25.86
Avg. acc. 97.02 89.60 / 68.99 66.03
GPT-4V Acc. with adv. pairs 85.57 72.13 47.82 29.85 18.49
Avg. acc. 92.03 85.32 / 65.20 57.18
Gemini Pro Acc. with adv. pairs 95.02 70.14 70.83 17.91 19.07
Avg. acc. 96.76 84.82 / 58.04 61.72